OpenStudy (anonymous):

csc(6x) + cot(6x)) / (csc(6x) - cot(6x)) = ________. 1) cot^2 (6x) 2) cot(3x) 3) cot^2(3x) 4) tan(6x)

OpenStudy (solomonzelman):

rafaeli , nominator = csc(6x) + cot(6x)) denominator = (csc(6x) - cot(6x)) correct?

OpenStudy (solomonzelman):

I see.

OpenStudy (anonymous):

yes that's correct!

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \sf \frac{Csc(x) +Cot(6x)}{Csc(6x) - Cot(6x)} }\) you can see that top and bottom are CONJUGATES ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so where do I go from that? :)

OpenStudy (solomonzelman):

Actually, I would simplifty top and bottom first. \(\Large\color{blue}{ \sf \frac{Csc(6x) +Cot(6x)}{Csc(6x) - Cot(6x)} }\) \(\Large\color{red}{ \sf Csc(6x)=\frac{1}{Sin(6x)} }\) \(\Large\color{red}{ \sf Cot(6x)=\frac{Cos(6x)}{Sin(6x)} }\) substitute..... \(\Huge\color{blue}{ \sf \frac{\frac{1}{Sin(6x)} +\frac{Cos(6x)}{Sin(6x)}}{\frac{1}{Sin(6x)} - \frac{Cos(6x)}{Sin(6x)}} }\) \(\Huge\color{blue}{ \sf \frac{\frac{1+Cos(6x)}{Sin(6x)} }{\frac{1-Cos(6x)}{Sin(6x)} } }\)

OpenStudy (anonymous):

Oooh I understood that. :) so how do i solve the rest now?

OpenStudy (anonymous):

do I flip the denominator fraction and multiply??

OpenStudy (solomonzelman):

YES !!!

OpenStudy (anonymous):

so i got 1+cos(6x) / 1-cos(6x)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{ \frac{ 1+Cos(x) }{ Sin(x) } }{ \frac{ 1-Cos(x) }{ Sin(x) } }}\) \(\Huge\color{blue}{ \sf \frac{ 1+Cos(x) }{ Sin(x) } \div \frac{ 1-Cos(x) }{ Sin(x) } =}\) \(\Huge\color{blue}{ \sf \frac{ 1+Cos(x) }{ Sin(x) } \times \frac{ Sin(x) }{ 1-Cos(x) } }\)

OpenStudy (solomonzelman):

@suhoping you had one mistake, you flip the fraction ....

OpenStudy (anonymous):

i thought i flipped it haha sorry

OpenStudy (solomonzelman):

it's alright :)

OpenStudy (anonymous):

the way you wrote it, you can just simplify the 6 out?

OpenStudy (solomonzelman):

Well ... it's really sin(6x). Last year, when I was doing calculations I was just writing Sin/Cos without saying Sin OF X and Cos OF X..

OpenStudy (solomonzelman):

Anyway, can you finish the problem off?

OpenStudy (anonymous):

oooh okay so I got this: 1+cos(6x) * 1-cos(6x)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{1+Cos(6x)}{Sin(6x)} \times \frac{Sin(6x)}{1-Cos(6x)} }\) \(\Huge\color{blue}{ \sf \frac{1+Cos(6x)}{\cancel{ Sin(6x)} } \times \frac{\cancel{ Sin(6x)} }{1-Cos(6x)} }\) \(\Huge\color{blue}{ \sf \frac{1+Cos(6x)}{1-Cos(6x)} }\) CORRECT!

OpenStudy (anonymous):

so wouldn't the answer really just be 1?

OpenStudy (solomonzelman):

Now \(\Huge\color{blue}{ \sf \frac{(1+Cos~6x~~)\color{red} { (1+Cos~6x~~) } }{(1-Cos~6x~~)\color{red} { (1+Cos~6x~~) }} }\)

OpenStudy (solomonzelman):

No, -cos on the bottom and +cos on the top, they are not even

OpenStudy (solomonzelman):

I am multiplying tp and bottom times [1+ cos 6x ]

OpenStudy (anonymous):

so the top would be 1+2cos(6x) ?

OpenStudy (solomonzelman):

don't forget about the last part of (a+b)(a+b)=a^2+2ab+b^2

OpenStudy (solomonzelman):

1+2cos(6x)+ cos ^2 (6x)

OpenStudy (solomonzelman):

@suhoping what would you get on the bottom?

OpenStudy (anonymous):

okay i see what you did. give me a second to figure out the answer

OpenStudy (solomonzelman):

can you tell me what is the bottom please

OpenStudy (anonymous):

1 + cos^2 (6x) ?

OpenStudy (solomonzelman):

Minus, not plus \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (solomonzelman):

1^2-cos^2 (6x) = 1 -cos^2x = Sin^2 (6x)

OpenStudy (anonymous):

so 1-cos^2 (6x) ?

OpenStudy (solomonzelman):

yes, and that's equal to sin^2 (6x) becuase cos^2 (6x) + sin^2 (6x) =1 - cos^2(6x) -cos^2 (6x) sin^2 (6x)=1-cos^2(6x)

OpenStudy (anonymous):

oh okay i got that

OpenStudy (anonymous):

so what do i do with the sin^2 (6x) and the 1+2cos(6x)+cos^2(6x) ?

OpenStudy (anonymous):

i divide them right?

OpenStudy (solomonzelman):

I am thinking to divide top and bottom by sin(6x)

OpenStudy (solomonzelman):

I mean ^2(6x)

OpenStudy (anonymous):

cos divided by sin is cot right?

OpenStudy (solomonzelman):

 (yes, you are right)